Four hamburgers and three orders of fries cost 4.65$. Five hamburgers and two orders of fries cost $4.85. How much does one hamburger and one order of fries cost?
Thank you!

Answer by lwsshak3(11628)

Four hamburgers and three orders of fries cost 4.65$. Five hamburgers and two orders of fries cost $4.85. How much does one hamburger and one order of fries cost?
let h=cost 0f one hamburger
let f=cost of 1 order of fries
4h+3f=4.65
5h+2f=4.85
8h+6f=9.30 (mult 1st eq by 2
15h+6f=14.55 (mult 2nd eq by 3
subtract:
7h=5.25
h=.75
3f=4.65-4h=1.65
cost 0f one hamburger=$.75
let f=cost of 1 orderr of fries=$1.65